Can you explain how to solve this word problem?:
How much of a 25% acid solution should be mixed with 5 L of a 15% acid solution to obtain a 20% acid solution?

Answer by mananth(16946)

How much of a 25% acid solution should be mixed with 5 L of a 15% acid solution to obtain a 20% acid solution?
..
25% acid solution................. x L
15% acid solution..................5 L
20% acid solution..................(x+5) L
..
25x+15*5=20(x+5)
25x+75=20x+100
25x-20x=100-75
5x=25
/5
x= 5 L.......... 25% solution